OpenStudy (agentnao):

What is the directrix of the parabola given by the equation x = 7y2 + 126y + 559?

OpenStudy (anonymous):

directrix is line perpendicular to axis of symmetry right?

OpenStudy (agentnao):

Yep!

OpenStudy (agentnao):

And that should be 7y^2 by the way.

OpenStudy (anonymous):

lol yh i figured

OpenStudy (anonymous):

crap i cant remember how to do these probs, give me a minute or two

OpenStudy (anonymous):

do we just solve for y values intersecting the y axis

OpenStudy (anonymous):

wait so shouldnt question be what is equation of parabola cause u already have equation of directrix

OpenStudy (anonymous):

hey wher'd u go?

OpenStudy (agentnao):

Sorry, my internet quit for a second. And no, that's not the equation of the directrix, thats the equation of the parabola.

OpenStudy (anonymous):

for a second i thought i was talking to my self :P

OpenStudy (agentnao):

Haha, no no. I wouldn't do that to you :)

OpenStudy (anonymous):

umm but equation of parabola cannot be in for x=ay^2+by +c

OpenStudy (anonymous):

can u check ur question again

OpenStudy (agentnao):

It's...well, that is what the question asks, but it's...

OpenStudy (anonymous):

lol sorry i cant help u i dont know how to solve it this way

OpenStudy (agentnao):

I don't know how to explain this without showing you. I'm doing a section of conics in my online class and you CAN have a parabola with x = blahblahblah, but I just need it to be written in general form

OpenStudy (anonymous):

repost it and some smart kid might be able to do it

OpenStudy (agentnao):

like a(x - h)^2 + h

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

hold up

OpenStudy (agentnao):

Haha, okay

OpenStudy (anonymous):

crap u are right so sorry, i think give me some time

OpenStudy (anonymous):

i think answer is 1/28

OpenStudy (anonymous):

in x = 7y2 + 126y + 559 the a value is = 1/4p hence it equals 1/28

OpenStudy (anonymous):

if thats not the answer i dont know it

OpenStudy (agentnao):

Oh my gosh, you're brilliant :) Thank you!

OpenStudy (anonymous):

aww shuks, thankx =)

OpenStudy (agentnao):

Oh wait...is it -8 and 1/28 or -7 and 27/28?

OpenStudy (anonymous):

no its p=1/28

OpenStudy (anonymous):

the line travelling along the x axis

OpenStudy (anonymous):

hey u there i wanna know if its right?

OpenStudy (agentnao):

Well, I won't know if it's right until I turn it in, but i don't want to turn it in if it's not right. It'ls a big catch 22.

OpenStudy (anonymous):

so lonely);

OpenStudy (anonymous):

whats the question? the one at the very top?

OpenStudy (anonymous):

yh

OpenStudy (agentnao):

The 28 part makes sense, that's part of the answer choices, but I need to know the rest.

OpenStudy (agentnao):

What is the directrix of the parabola given by the equation x = 7y2 + 126y + 559?

OpenStudy (anonymous):

lol im only in grade 11 so most of this beyond me

OpenStudy (agentnao):

Lol, I'm in grade 11 too :)

OpenStudy (agentnao):

No worries though, I'm not understanding this question any better than you are and I'm looking at the course material right now!

OpenStudy (anonymous):

lol i never learnt this, i looked it up once for a contest math problem but never did it

OpenStudy (anonymous):

man my school sux i learn nothin

OpenStudy (anonymous):

im trying to find the formula for the directrix. google is failing me lol =/

OpenStudy (agentnao):

Again, no worries. My school is so terrible, I'm transferring to a new school in a whole 'nother city and I'm not moving there either. I'll just be driving 30 minutes each way.

OpenStudy (anonymous):

wikipedia it

OpenStudy (anonymous):

i would start by completing the square though, most likely that would be the first step to figuring out the directrix.

OpenStudy (agentnao):

Oh! Joe, I can help you with that one. The directrix is x = h - c

OpenStudy (agentnao):

Hahd figured out c, but now we need h

OpenStudy (anonymous):

lets complete the square first just to make the equation look nice: \[7y^2+126y+559 = 7(y^2+18y+81)+559-567 = 7(y+9)^2-8\]

OpenStudy (anonymous):

so now we have: \[x = 7(y+9)^2-8 \iff x+8 = 7(y+9)^2 \iff \frac{1}{7}(x+8) = (y+9)^2\]

OpenStudy (agentnao):

oh my...

OpenStudy (agentnao):

Holy cow! That's super close to the form we need to get it in!

OpenStudy (anonymous):

according to some websites, the standard form of a parabola is: \[(y-k)^2 = 4p(x-h)\]

OpenStudy (anonymous):

the 4p is going to equal that fraction we have, so: \[4p = \frac{1}{7} \iff p = \frac{1}{28}\] The directrix is going to be: \[x = -p+h = -\frac{1}{28}+ (-8) = -\frac{225}{28} \]

OpenStudy (agentnao):

<3

OpenStudy (agentnao):

You're awesome. Thanks so much!

OpenStudy (anonymous):

man my computer died

OpenStudy (anonymous):

i had it too 0;

OpenStudy (anonymous):

i did the right thing calling joemath

OpenStudy (agentnao):

You called him?

OpenStudy (anonymous):

its hard to remember all these parabola formulas >.<

OpenStudy (anonymous):

yh

OpenStudy (anonymous):

he's helped me before

OpenStudy (anonymous):

plus he's in uni so its like help from an older friend

OpenStudy (agentnao):

How do you call someone on here? Or do you know him personally? And YES. It is extremely difficult. The amount of formulas there are for arithmetic and geometric sequences alone is enough to kill you.

OpenStudy (anonymous):

lol when u see them answering someones answer spam that wall

OpenStudy (agentnao):

Thanks to both of you though. You were incredibly helpful.

OpenStudy (anonymous):

meh joemath is awesome

OpenStudy (agentnao):

Hahahaha, I will certainly keep that in mind for future references :)

OpenStudy (agentnao):

Yes he is!

OpenStudy (anonymous):

psh im awesome too just at stuff i know bye

OpenStudy (agentnao):

Yes, you were definitely awesome too :) Farewell!

OpenStudy (anonymous):

hahd is very awesome :)

OpenStudy (anonymous):

LOL thnkx

OpenStudy (anonymous):

the good thing is i actually learned stuff so i feel good

OpenStudy (anonymous):

thats why i stay around lol. sometimes you learn a new technique to solve something, or maybe you learn to solve something you couldnt before.